# Trajectory Planning for Autonomous Vehicle Using Iterative Reward Prediction in Reinforcement Learning

### v3

@@ Abstract
Traditional trajectory planning methods for autonomous vehicles have several limitations. For example, heuristic and explicit simple rules limit generalizability and hinder complex motions. These limitations can be addressed using reinforcement learning-based trajectory planning. However, reinforcement learning suffers from unstable learning and existing reinforcement learning-based trajectory planning methods do not consider the uncertainties. Thus, this paper, proposes a reinforcement learning-based trajectory planning method for autonomous vehicles. The proposed method involves an iterative reward prediction approach that iteratively predicts expectations of future states. These predicted states are then used to forecast rewards and integrated into the learning process to enhance stability. Additionally, a method is proposed that utilizes uncertainty propagation to make the reinforcement learning agent aware of uncertainties. The proposed method was evaluated using the CARLA simulator. Compared to the baseline methods, the proposed method reduced the collision rate by 60.17%, and increased the average reward by 30.82 times. A video of the proposed method is available at https://www.youtube.com/watch?v=PfDbaeLfcN4.
@@ I INTRODUCTION
Since the 2007 DARPA Urban Challenge [1], autonomous vehicles(AVs) have been studied intensively. In additon, planning methods for AVs have also been researched intensively and the findings have allowed AVs to drive successfully within limited areas, using rule-based and optimization-based algorithms, explicit heuristic rules, and parameters specified for the given area. However, these traditional approaches suffer from several limitations, including a lack of generality and a lack of complex motion. For example, the heuristic rules and parameters specified for the given area may not be applied to the other areas which impacts the scalability. In addition, many possible scenarios must be considered in real-world applications. If various scenarios are generalized with few scenarios(e.g., lane following, and lane changing), an overly simple policy could be obtained. Numerous studies have employed deep learning to address these limitations [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
The most popular approach is imitation learning(IL), which learns a driving policy directly from expert driving data [2, 3]; however, IL also has several limitations, including the cost of scalability, simple driving policies, and safety. For example to scale up the AV using the IL method, expert data must be obtained for all scenarios and targeted areas, which is costly. In terms of driving policies, IL is typically used for handling simpler driving tasks, e.g., lane following. To learn complex policies or policies in corner cases, it should have a lot of data which can be costly and time-consuming. In addition, an expert’s demonstrated policies for complex scenarios or corner cases are more distributed than the simple scenarios, which may yield large errors or learning may be infeasible [4]. Relative to safety limitations, insufficient amounts of data on dangerous cases or corner cases are available; thus, the IL agent could output dangerous policies due to a lack of training data.
Figure 1: The AV(red car) plans a trajectory(red boxes) along the lane while avoiding a parked car(blue car) and a car that is changing lanes(blue car) from the left lane. Other vehicles’ predicted trajectories are represented as blue boxes and trajectories considering uncertainty are represented as rounded light blue boxes. The goal of each AV’s trajectory is determined iteratively in each time step from the previously predicted state of the AV and the states of other vehicles. The AV’s states and corresponding goals are represented as red, orange, yellow, green, blue, and purple circles(shown in chronological order). The AV’s planned trajectories are shown as red boxes and trajectories considering uncertainty are shown as rounded light red boxes. The predicted states of other vehicles and the iteratively predicted states of the AV are employed for prediction of reward within the planning horizon which makes the RL learning process stable.
@@
Thus this study, proposes an RL-based trajectory planning method for AVs that overcomes the identified limitations of RL and traditional planning methods. The proposed method employs a reward prediction(RP) during the learning process which predicts expectations of future states. These predicted states are then used to forecast rewards and integrated into the learning process to enhance stability. In addition, an iterative RP(IRP) method that uses RP iteratively to predicts relevant states, actions, and corresponding rewards accurately is employed. As a result, the performance of the agent and the learning stability are increased. Additionally, a method is proposed that utilizes uncertainty propagation [21, 22] to make the reinforcement learning agent aware of uncertainties. An overview of the proposed method is shown in Fig.1.
@@ III Method
To solve trajectory planning problems using RL, RL algorithms with continuous action space, e.g., the DDPG [26], TD3 [20], and PPO [27] algorithms, are more suitable than algorithms that utilize a discrete action space. Since getting smooth behaviors requires an increase in the size of the discrete action space which leads to discrete control methods being intractable. In addition, for simplicity and interpretability, algorithms with deterministic policies(rather than stochastic policies) are selected. Furthermore, generating the goal for a trajectory is a better action choice than control command because more specific policies are required for control command methods to yield rewards successfully. Thus, the proposed trajectory planning method generates a goal using a deterministic continuous control RL algorithm, and we assume that information about the localization, route path, trajectory prediction of other traffic participants, and object detection is given.
@@ III-A Problem Formulation
We formulate the trajectory planning problem as a Markov decision process, which is defined by the tuple $(S,A,P,R)$. Here, $s\in S$ is the continuous state space, $a\in A$ is the continuous action space, $P$ is the probability of state transition, and $R$ is the reward received after each transition. The return is defined as discounted sum of rewards $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}$, where $\gamma\in(0,1)$ is the discount factor determining the priority of short-term rewards. The purpose of RL is to learn an optimal policy that maximizes the expected accumulated rewards as follows:
@@
where $r$ is the reward function, $\rho^{\pi}$ is the state distribution under the policy $\pi$.
@@ III-B Reward Prediction
The objective function $J(\pi_{\theta})$ in (2) can be defined as the action value function $Q$, the advantage function $A$, and the TD error $\delta$. The action value and advantage functions are approximated with neural networks in deep RL method; however, function approximation with neural network always involves errors, which causes instability during the learning process and poor performance. To improve learning stability, we utilize N-step SARSA concept. N-step SARSA improves the learning stability by utilizing the error reduction property of n-step returns. However, the rewards utilized in N-step SARSA includes randomness and variance due to the randomness of state transitions, which also causes learning instability.
We propose an RP method that utilizes the error reduction property of the n-step returns and resolves the aforementioned problem caused by variance of rewards. Rewards are predicted by predicting the expectations of future states. Unlike N-step SARSA, the expected return from the action value function of RP is not based on average rewards resulting from state transitions, but rather rewards associated with expected states. Thus, we exploit the advantages of N-step SARSA while reducing the effect by the variance of the rewards. The AV framework involves the prediction of traffic participants and the planned trajectory of the AV can be obtained using the output goal of the RL agent. Thus, future rewards can be predicted by utilizing the planned future states of the AV and the predicted states of other traffic participants. The proposed RP method utilizes the Bellman equation and is employed during the action value function update process as follows:
@@
$\begin{split}&s^{e}_{t,\tau}\in\mathcal{T}^{t}_{e},\;s^{o,k}_{t,\tau}\in\mathcal{T}^{t}_{o,k},\;s^{o}_{t,\tau}=\{s^{o,}_{t,\tau},s^{o,1}_{t,\tau},\cdots s^{o,n}_{t,\tau}\}\;\\ &s_{t,\tau}=f(s^{e}_{t,\tau},s^{o}_{t,\tau}),\;r_{t,\tau+\delta}=g(s_{t,\tau},s_{t,\tau+\delta}),\\ &J(\pi)=Q^{\pi}(s_{t},a_{t}),\\ &L(\theta^{Q})=\mathbb{E}_{s_{t}\sim\rho^{\pi},a_{t}\sim\pi_{\theta}}[(Q^{\pi}(s_{t},a_{t})-y_{t})^{2}],\\ &y_{t}=r_{t,\delta}+\gamma\cdot r_{t,2\delta}+\cdots\\ &\quad\quad\quad+\gamma^{T/\delta-1}\cdot r_{t,T}+\gamma^{T/\delta}\cdot Q^{\pi}(s_{t,T},\pi(s_{t,T})),\\ \end{split}$ (4)
@@
where $T$ is a planning/prediction horizon, $\tau$ is the time within the planning horizon $\tau\in[0,T]$, $s^{e}_{t,\tau}$ is the predicted future state of the AV at time $t+\tau$ from trajectory $\mathcal{T}^{e}_{t}$ planned at time $t$ using the goal $\pi(s_{t,0})$, $s^{o,k}_{t,\tau}$ is the predicted future state of the $k$th other traffic participant at time $t+\tau$ from prediction $\mathcal{T}^{o,k}_{t}$ predicted at time $t$, $s^{o}_{t,\tau}$ is the predicted future states of $n$ other traffic participants at time $t+\tau$ from prediction at time $t$. In addition, $f$ is a function that outputs the state $s_{t,\tau}$ for the RL agent by combining the predicted future states of the AV $s^{e}_{t,\tau}$ and the other traffic participants $s^{o}_{t,\tau}$, $\delta$ is a sufficiently small time step size, $g$ is a function that predicts the reward $r_{t,\tau+\delta}$ at time $t+\tau+\delta$ during transition from $s_{t,\tau}$ to $s_{t,{\tau+\delta}}$ , $L(\theta^{Q})$ is the loss for the action value function $Q^{\pi}$, and $y_{t}$ is the expected return inferred by utilizing predicted rewards and the action value of the predicted state at the planning horizon, and its corresponding action. Fig.2 shows the predicted states of the other vehicle and the AV with its goal at time $t$.
Although the RP stabilizes learning by utilizing the expectations of future states, its variance is not considered; however, in terms of predicting the states of other traffic participants and the AV, its variance is strongly related to safety. Thus, we utilize the uncertainty propagation on future states of other traffic participants and the AV to consider the variance. Additional details about the uncertainty propagation process are given in Section III-D.
@@
Figure 3: Demonstration of inaccurate RP. Here, the agent has planned a trajectory that makes the AV to be close to the reference. The trajectories planned at each time step beginning from each state action pair are represented by different colors. The return predicted by RP, $Q_{RP}$, is based on the red trajectory planned at time $t$. The agent plans a new trajectory at each step; thus, the planned trajectory that RP utilized at time $t$(red) and actual trajectory executed by the AV(blue) are different. Thus, $Q_{RP}$ has an error of $\epsilon$ between the true return $Q_{true}$ which is a return obtained by following the blue trajectory.
@@ III-C Iterative Reward Prediction
Note that the rewards predicted by the RP still is inaccurate and has errors from the true rewards. Even though the AV can track the given trajectory perfectly, the planned trajectory in the next time step will differ slightly from the previously planned trajectory, which causes inaccurate prediction of reward. In turn, inaccurate prediction of reward results in learning instability and poor performance. Fig.3 demonstrates the inaccuracy problem of RP.
@@
To resolve the inaccuracy problem of RP, we propose the IRP method, which predicts the reward by iteratively planning a new trajectory at the predicted state and predicting the reward of that trajectory. Fig. 4 shows how the proposed IRP method operates. Compared to the conventional RP, the proposed IRP predicts the reward more accurately and reduces the error of the function approximation of the action value function $Q^{\pi}$. The mathematical representation of the IRP is given as follows:
@@
$\begin{split}&s_{t+\tau,\delta}=h(s_{t+\tau,0},\pi(s_{t+\tau,0})),\\ &s_{t+\tau+\delta,0}=s_{t+\tau,\delta},\\ &r_{t+\tau,\delta}=g(s_{t+\tau,0},s_{t+\tau,\delta}),\\ &J(\pi)=Q^{\pi}(s_{t},a_{t}),\\ &L(\theta^{Q})=\mathbb{E}_{s_{t}\sim\rho^{\pi},a_{t}\sim\pi}[(Q^{\pi}(s_{t},a_{t})-y_{t})^{2}],\\ &y_{t}=r_{t,\delta}+\gamma\cdot r_{t+\delta,\delta}+\cdots+\gamma^{T/\delta-1}\cdot r_{t+T-\delta,\delta}\\ &\quad\quad\quad+\gamma^{T/\delta}\cdot Q^{\pi}(s_{t+T-\delta,\delta},\pi(s_{t+T-\delta,\delta})),\\ \end{split}$ (5)
@@
where time $t$ is when planning first started, $s_{t+\tau,\delta}$ is the predicted state at time $t+\tau+\delta$ predicted from the previously predicted state at time $t+\tau$, and $h$ is a function that outputs a state $s_{t+\tau,\delta}$. Here, the predicted states $s_{t+\tau,\delta}=s_{t+\tau+\delta,0}$ are used iteratively to predict the next states $s_{t+\tau+\delta,\delta}$. In function $h$, the future states of the other traffic participants are assumed to follow the trajectories predicted at time $t$, which is assumed to be given. However, the future states of the AV are updated iteratively at each time step.
@@
In addition, considering the agent’s future action more accurately enables the agent to drive along a complicated trajectory and stabilizes the learning process because the agent has more comprehensive understanding of its own future actions. As shown in Fig.4, the AV can plan a red trajectory by considering agent’s future actions; thus, it is capable of driving a complicated trajectory.
@@
Figure 4: Predicted states(blue boxes) of other traffic participants(blue cars) and predicted states(red boxes) of the AV(red car) and actions of the RL agent. The action of the RL agent is determined iteratively from the previous state of the AV, its goal, and the states of the other vehicles. The goals and the AV’s start state are represented as red, orange, yellow, green, blue, and purple circles in chronological order. The predicted states of the other vehicles and the iteratively predicted states of the AV are utilized to predict the reward during the learning process.

### v4

@@ Abstract
Traditional trajectory planning methods for autonomous vehicles have several limitations. For example, heuristic and explicit simple rules limit generalizability and hinder complex motions. These limitations can be addressed using reinforcement learning-based trajectory planning. However, reinforcement learning suffers from unstable learning, and existing reinforcement learning-based trajectory planning methods do not consider the uncertainties. Thus, this paper, proposes a reinforcement learning-based trajectory planning method for autonomous vehicles. The proposed method involves an iterative reward prediction approach that iteratively predicts expectations of future states. These predicted states are then used to forecast rewards and integrated into the learning process to enhance stability. Additionally, a method is proposed that utilizes uncertainty propagation to make the reinforcement learning agent aware of uncertainties. The proposed method was evaluated using the CARLA simulator. Compared to the baseline methods, the proposed method reduced the collision rate by 60.17%, and increased the average reward by 30.82 times. A video of the proposed method is available at https://www.youtube.com/watch?v=PfDbaeLfcN4.
@@ I INTRODUCTION
Since the 2007 DARPA Urban Challenge [1], autonomous vehicles(AVs) have been studied intensively. In addition, planning methods for AVs have also been researched intensively and the findings have allowed AVs to drive successfully within limited areas, using rule-based and optimization-based algorithms, explicit heuristic rules, and parameters specified for the given area. However, these traditional approaches suffer from several limitations, including a lack of generality and a lack of complex motion. For example, the heuristic rules and parameters specified for the given area may not be applied to the other areas, which impact the scalability. In addition, many possible scenarios must be considered in real-world applications. If various scenarios are generalized with few scenarios(e.g., lane following, and lane changing), an overly simple policy could be obtained. Numerous studies have employed deep learning to address these limitations [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
The most popular approach is imitation learning(IL), which learns a driving policy directly from expert driving data [2, 3]; however, IL also has several limitations, including the cost of scalability, simple driving policies, and safety. For example, to scale up the AV using the IL method, expert data must be obtained for all scenarios and targeted areas, which is costly. In terms of driving policies, IL is typically used for handling simpler driving tasks, e.g., lane following. To learn complex policies or policies in corner cases, it should have a lot of data which can be costly and time-consuming. In addition, an expert’s demonstrated policies for complex scenarios or corner cases are more distributed than the simple scenarios, which may yield large errors or learning may be infeasible [4]. Relative to safety limitations, insufficient amounts of data on dangerous cases or corner cases are available; thus, the IL agent could output dangerous policies due to a lack of training data.
Figure 1: The AV(red car) plans a trajectory(red boxes) along its lane while avoiding a parked car(blue car) on the right and a car that is changing lanes(another blue car) from the left lane. The AV’s planned trajectories are represented as red boxes, and their trajectories considering uncertainty, are represented as rounded light red boxes. The other vehicles’ predicted trajectories are represented as blue boxes, and their trajectories, considering uncertainty, are represented as rounded light blue boxes. The goal of each AV’s trajectory is determined iteratively at each time step from the previously predicted state of the AV and the states of the other vehicles. The AV’s states and corresponding goals are represented by red, orange, yellow, green, blue, and purple circles(shown in chronological order). The predicted states of the other vehicles and the iteratively predicted states of the AV are employed for the prediction of reward within the planning horizon, which stabilizes the RL learning process.
@@
Thus, this study proposes an RL-based trajectory planning method for AVs that overcomes the identified limitations of RL and traditional planning methods. The proposed method employs a reward prediction(RP) during the learning process which predicts expectations of future states. These predicted states are then used to forecast rewards and integrated into the learning process to enhance stability. In addition, an iterative RP(IRP) method that uses RP iteratively to predicts relevant states, actions, and corresponding rewards accurately is employed. As a result, the performance of the agent and the learning stability are increased. Additionally, a method is proposed that utilizes uncertainty propagation [21, 22] to make the reinforcement learning agent aware of uncertainties. An overview of the proposed method is shown in Fig.1.
@@ III Method
To solve trajectory planning problems using RL, RL algorithms with continuous action space, e.g., the DDPG [26], TD3 [20], and PPO [27] algorithms, are more suitable than algorithms that utilize a discrete action space. Since getting smooth behaviors requires an increase in the size of the discrete action space which leads to discrete control methods being intractable. In addition, for simplicity and interpretability, algorithms with deterministic policies(rather than stochastic policies) are selected. Furthermore, generating the goal for a trajectory is a better action choice than control command because more specific policies are required for control command methods to yield rewards successfully. Thus, the proposed trajectory planning method generates a goal using a deterministic continuous control RL algorithm. It is assumed that information about the localization, route path, trajectory prediction of other traffic participants, and object detection is provided. However, trajectory prediction of other traffic participants is only used during the learning process.
@@ III-A Problem Formulation
The trajectory planning problem is formulated as a Markov decision process, which is defined by the tuple $(S,A,P,R)$. Here, $s\in S$ is the continuous state space, $a\in A$ is the continuous action space, $P$ is the probability of state transition, and $R$ is the reward received after each transition. The return is defined as discounted sum of rewards $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}$, where $\gamma\in(0,1)$ is the discount factor determining the priority of short-term rewards. The purpose of RL is to learn an optimal policy that maximizes the expected cumulative reward as follows:
@@
where $r$ is the reward function, $\rho^{\pi}$ is the state distribution under the policy $\pi$. The future states of the AV and other traffic participants are assumed to follow the planned trajectory and the predicted trajectory, respectively, with deviations following a normal distribution, as assumed in [21].
@@ III-B Reward Prediction
The objective function $J(\pi_{\theta})$ in (2) can be defined as the action value function $Q$, the advantage function $A$, and the TD error $\delta$. The action value and advantage functions are approximated with neural networks in the deep RL method; however, function approximation with neural network always involves errors, which causes instability during the learning process and poor performance. To address this, the N-step SARSA concept is employed in RP. N-step SARSA improves learning stability by utilizing the error reduction property of n-step returns. Nevertheless, the learning process of N-step SARSA remains susceptible to instability.
To overcome this, the RP method that utilizes the error reduction property of the n-step returns and enhances the learning stability is proposed. To achieve this stability, RP enables the agent to explicitly learn the consequences of actions. This is accomplished by utilizing predictions of traffic participants’ behavior and the planned trajectory of the AV, which is obtained using the agent’s output goal. Specifically, future rewards are predicted by considering the planned future states of the AV and the predicted states of other traffic participants. Thus, RP leverages the advantages of N-step SARSA while explicitly learning action consequences. These two traits of RP contribute to stabilizing the learning process. The proposed RP method is employed during the action value function update process as follows:
@@
$\begin{split}&s^{e}_{t,\tau}\in\mathcal{T}^{e}_{t},\;s^{o,k}_{t,\tau}\in\mathcal{T}^{o,k}_{t},\;s^{o}_{t,\tau}=\{s^{o,}_{t,\tau},s^{o,1}_{t,\tau},\cdots s^{o,n}_{t,\tau}\}\;\\ &s_{t,\tau}=f(s^{e}_{t,\tau},s^{o}_{t,\tau}),\;R_{t,\tau+\delta}=g(s_{t,\tau},s_{t,\tau+\delta}),\\ &J(\pi)=Q^{\pi}(s_{t},a_{t}),\\ &L(\theta^{Q})=\mathbb{E}_{s_{t}\sim\rho^{\pi},a_{t}\sim\pi_{\theta}}[(Q^{\pi}(s_{t},a_{t})-y_{t})^{2}],\\ &y_{t}=R_{t,\delta}+\gamma\cdot R_{t,2\delta}+\cdots\\ &\quad\quad\quad+\gamma^{T/\delta-1}\cdot R_{t,T}+\gamma^{T/\delta}\cdot Q^{\pi}(s_{t,T},\pi(s_{t,T})),\\ \end{split}$ (4)
@@
where time $t$ is when the AV’s trajectory is planned and the trajectory prediction of other traffic participants is made, $T$ is a planning/prediction horizon, $\delta$ is a sufficiently small time step size, $\tau$ is the time within the planning horizon $\tau\in[\delta,T]$, $s^{e}_{t,\tau}$ is the predicted future state of the AV at time $t+\tau$ from trajectory $\mathcal{T}^{e}_{t}$ planned at time $t$ using the goal $\pi(s_{t,0})$, $s^{o,k}_{t,\tau}$ is the predicted future state of the $k$th other traffic participant at time $t+\tau$ from prediction $\mathcal{T}^{o,k}_{t}$ predicted at time $t$, $s^{o}_{t,\tau}$ is the predicted future states of $n$ other traffic participants at time $t+\tau$ from prediction at time $t$. In addition, $f$ is a function that combines the predicted future states of the AV and the other traffic participants at the same time to output the state $s_{t,\tau}$ for the RL agent, $g$ is a function that predicts the reward $R_{t,\tau+\delta}$ at time $t+\tau+\delta$ during transition from $s_{t,\tau}$ to $s_{t,\tau+\delta}$, $L(\theta^{Q})$ is the loss for the action value function $Q^{\pi}$, $y_{t}$ is the expected return inferred by utilizing predicted rewards and the action value of the predicted state at the planning horizon, and its corresponding action, and $R_{t,\delta}$ is a reward received from the environment. Predicted future states are used to predict rewards, which are then utilized during the action value function update process. Fig.2 shows the predicted states of the other vehicle and the AV with its goal at time $t$.
Unlike N-step SARSA, the expected return from the action value function learned by RP is not based on rewards resulting directly from state transitions, but rather on rewards associated with expected states. Namely, RP increases learning stability by utilizing expected future states, yet it does not directly account for the variation of these future states. However, in terms of predicting the states of other traffic participants and the AV, its variance is strongly related to safety. Therefore, uncertainty propagation on future states of other traffic participants and the AV is utilized to consider this variance. Additional details about the uncertainty propagation process are given in Section III-D.
@@
Figure 3: Demonstration of inaccuracy problem of RP. Here, the agent has planned a trajectory that takes AV close to the reference. The trajectories planned at each time step beginning from each state-action pair are represented by different colors. The return predicted using RP, $Q_{RP}$, is based on the red trajectory planned at time $t$. The agent plans a new trajectory at each step; therefore, the planned trajectory that RP utilized at time $t$(red) and the actual trajectory executed by the AV(blue) are different. Thus, $Q_{RP}$ has an error $\epsilon$ compared to the true return $Q_{true}$, which is the return obtained by following the blue trajectory.
@@ III-C Iterative Reward Prediction
In general, AVs plan a trajectory at every time step, which is a common practice due to potential inaccuracies in predicting the behavior of other traffic participants and the imprecise path tracking by the controller. Therefore, the proposed method also assumes trajectory planning at every time step. Given this assumption, even if the controller tracks the given trajectory perfectly, there is no guarantee that the trajectory generated using the agent’s output will be the same as the previous one. Consequently, this inaccuracy of RP leads to inaccurate prediction of rewards, resulting in learning instability and poor performance. Fig.3 demonstrates the inaccuracy problem of RP.
@@
To address the inaccuracy problem of RP, the IRP method is proposed. This method predicts the rewards by iteratively planning a new trajectory at the predicted state and predicting the reward of that trajectory for one time step. Fig. 4 shows how the proposed IRP method operates. Compared to the conventional RP, the proposed IRP predicts the rewards more accurately and reduces the error of the function approximation of the action value function $Q^{\pi}$. The mathematical representation of the IRP is given as follows:
@@
$\begin{split}&s_{t+\tau,0}=f(s^{e}_{t+\tau,0},s^{o}_{t,\tau})\\ &s_{t+\tau,\delta}=h(s_{t+\tau,0},\pi(s_{t+\tau,0})),\\ &R_{t+\tau,\delta}=g(s_{t+\tau,0},s_{t+\tau,\delta}),\\ &J(\pi)=Q^{\pi}(s_{t},a_{t}),\\ &L(\theta^{Q})=\mathbb{E}_{s_{t}\sim\rho^{\pi},a_{t}\sim\pi}[(Q^{\pi}(s_{t},a_{t})-y_{t})^{2}],\\ &y_{t}=R_{t,\delta}+\gamma\cdot R_{t+\delta,\delta}+\cdots+\gamma^{T/\delta-1}\cdot R_{t+T-\delta,\delta}\\ &\quad\quad\quad+\gamma^{T/\delta}\cdot Q^{\pi}(s_{t+T-\delta,\delta},\pi(s_{t+T-\delta,\delta})),\\ \end{split}$ (5)
@@
where time $t$ is when planning first started, $s_{t+\tau,\delta}$ is the predicted state at time $t+\tau+\delta$ predicted from the previously predicted state at time $t+\tau$, and $h$ is a function that outputs a predicted state using the current state, and action. Here, the predicted state, such as $s_{t+\tau,\delta}$ is used as $s_{t+\tau+\delta,0}$ for iteratively predicting the next state $s_{t+\tau+\delta,\delta}$. In function $h$, the future states of the other traffic participants are assumed to follow the trajectories predicted at time $t$, which is assumed to be given. However, the future states of the AV are updated iteratively at each time step.
@@
In addition, considering the agent’s future action more accurately enables the agent to drive along a complicated trajectory because the agent has a more comprehensive understanding of its future actions. As shown in Fig.4, the AV can plan a red trajectory by considering the agent’s future actions; thus, it is capable of driving a complicated trajectory.
@@
Figure 4: Predicted states(blue boxes) of other traffic participants(blue cars) and predicted states(red boxes) of the AV(red car) and actions of the RL agent. The action of the RL agent is determined iteratively from the previous state of the AV, its goal, and the states of the other vehicles. The goals and the AV’s start states are represented as red, orange, yellow, green, blue, and purple circles in chronological order. The predicted states of the other vehicles and the iteratively predicted states of the AV are utilized to predict the rewards during the learning process.